### v2

# MBBQ: A Dataset for Cross-Lingual Comparison of Stereotypes in Generative LLMs

Abstract
Generative large language models (LLMs) have been shown to exhibit harmful biases and stereotypes. While safety fine-tuning typically takes place in English, if at all, these models are being used by speakers of many different languages. There is existing evidence that the performance of these models is inconsistent across languages and that they discriminate based on demographic factors of the user. Motivated by this, we investigate whether the social stereotypes exhibited by LLMs differ as a function of the language used to prompt them, while controlling for cultural differences and task accuracy. To this end, we present MBBQ (Multilingual Bias Benchmark for Question-answering), a carefully curated version of the English BBQ dataset extended to Dutch, Spanish, and Turkish, which measures stereotypes commonly held across these languages. We further complement MBBQ with a parallel control dataset to measure task performance on the question-answering task independently of bias. Our results based on several open-source and proprietary LLMs confirm that some non-English languages suffer from bias more than English, even when controlling for cultural shifts. Moreover, we observe significant cross-lingual differences in bias behaviour for all except the most accurate models. With the release of MBBQ, we hope to encourage further research on bias in multilingual settings. The dataset and code are available at https://github.com/Veranep/MBBQ.
1 Introduction
Generative large language models (LLMs) have proven useful for tasks ranging from summarization, translation and writing code to answering healthcare and legal questions and taking part in open-domain dialogue (Bang et al., 2023; Zan et al., 2023; Hung et al., 2023). At the same time, a large amount of work has shown that they exhibit various harmful biases and stereotypes (e.g., Dinan et al., 2020; Esiobu et al., 2023; Cheng et al., 2023; Jeoung et al., 2023; Plaza-del Arco et al., 2024), and engage with harmful instructions (Zhang et al., 2023). Yet, LLMs are being used by vast amounts of speakers over the world. Although most models have not intentionally and systematically been trained to be multilingual—with English being the overwhelmingly dominant language in the training data—they are actively being used by speakers of at least 150 different languages (Zheng et al., 2024). However, if they have received any sort of safety training, this is often only in English (Touvron et al., 2023). Given this, combined with evidence that LLMs show differences in performance across languages (Holtermann et al., 2024) and can be inconsistent cross-linguistically when asked about factual knowledge (Ohmer et al., 2023; Qi et al., 2023), we hypothesise that the social biases exhibited by an LLM may differ as a function of the language used to prompt it. We believe that shedding light on this issue is critical to arrive at a more comprehensive overview of bias in NLP and ultimately improve model fairness.
To make progress in this direction, in this paper we investigate to what extent the presence of bias regarding social stereotypes differs when chat-optimised generative LLMs are prompted in different languages, while controlling for cultural idiosyncrasies and model accuracy.
Models have been shown to display different kinds of biases depending on how users describe themselves demographically (Smith et al., 2022) and to discriminate against speakers of African American English when making decisions about character or criminality (Hofmann et al., 2024). Thus, the language employed by a user to prompt a model could cause models to generate responses that exhibit varied harmful properties, possibly due to different languages being underrepresented to different degrees in the training data. While bias benchmarks for non-English languages are challenging to develop and hence rare (Talat et al., 2022), several recent studies investigate generative LLM’s safety and biases in languages other than English (Zhang et al., 2023; Shen et al., 2024; Zhao et al., 2024). However, they either investigate one particular bias, or consider the safety or bias of a model as a whole without investigating the exact biases present; do not control for cross-linguistic differences in task performance; and do not focus on the comparison of model biases across languages. To address these gaps, we adapt the approach to bias evaluation for multiple stereotype categories proposed by Parrish et al. (2022), who originally evaluated English-based question answering models, to the conversational, generative setting and extend it to three additional languages.
Concretely, we translate the Bias Benchmark for Question-answering (BBQ; Parrish et al., 2022) from English into Dutch, Spanish, and Turkish. This dataset consists of multiple-choice questions referring to stereotypes from a wide variety of social categories, including age, socioeconomic status, and gender identity, among others. We carry out a careful manual analysis to retain only those stereotypes that are common to the four languages we consider. This contrasts with the approach of Jin et al. (2023), who have recently translated BBQ into Korean and extended it to adapt it to the South Korean cultural context. While capturing and accounting for cultural differences is an important challenge (Talat et al., 2022; Arora et al., 2023), our aim here is to investigate whether models behave differently across languages regarding common stereotypes. To our knowledge, we are the first to investigate this question in generative LLMs. In addition, in order to separate a model’s performance on the question-answering task from the measured biases, we devise a parallel control set. We require this control set to measure task performance independent from bias, because only if models show similar task performance across languages can we attribute measured differences in bias scores to biased model behavior (Levy et al., 2023).
In summary, our main contributions are: (1) We present the Multilingual Bias Benchmark for Question-answering (MBBQ), a hand-checked translation of the English BBQ dataset into Dutch, Spanish, and Turkish for measuring cross-lingual differences on a subset of stereotypes widely held in these languages. (2) We create a parallel MBBQ control dataset to test for task performance independently from bias; both MBBQ and its control counterpart will be publicly released to facilitate further research as well as possible dataset extensions (to other languages and/or stereotypes) in the future. (3) We carry out experiments with $7$ LLMs comparing accuracy on the question-answering task and bias behaviour across $6$ bias categories in the $4$ languages mentioned above. Our results show that all models display significant differences across languages in question-answering accuracy and, with the exception of the most accurate models, also in bias behavior—despite controlling for cultural shifts. When bias scores differ significantly across languages, models are generally most biased in Spanish, and least biased in English or Turkish. Models are generally less accurate and give more biased answers when the context of a question is ambiguous, relying on stereotypes rather than acknowledging that the question cannot be answered.
Overall, our findings highlight the importance of controlling for cultural differences and task accuracy when measuring model bias. With MBBQ, we hope to encourage further work on bias in multilingual settings and facilitate research on cross-lingual debiasing.
2 Related work
Social biases in NLP
There is a considerable number of works that detect, evaluate and mitigate social biases in NLP; see Dev et al. (2022) and Gallegos et al. (2023) for comprehensive overviews of harms present in NLP technologies and existing ways to measure them. Earlier work on static word embeddings often compared words of interest, e.g., profession terms, to word lists that capture two demographic groups, for instance men and women for gender bias (Bolukbasi et al., 2016; Caliskan et al., 2017). If a word of interest is more similar to one word list than to the other, this reflects a bias in the corresponding word embedding. More recently, research on bias in language models has uncovered a wide range of biases present in those models, typically through bias measures defined on a specific benchmark dataset (Dev et al., 2022; van der Wal et al., 2024). The StereoSet (Nadeem et al., 2021) and CrowS-pairs (Nangia et al., 2020) datasets have identified biases about demographic groups associated with attributes such as gender, race, and nationality in (masked) language modeling, and similar datasets exist for many downstream tasks (Dev et al., 2022), including question answering (Li et al., 2020; Parrish et al., 2022) and dialogue generation (Dinan et al., 2020; Liu et al., 2020a; b). [^1] See https://github.com/i-gallegos/Fair-LLM-Benchmark for a list of bias evaluation datasets.
In a non-English setting, Névéol et al. (2022) have translated the CrowS-Pairs dataset to French. Reusens et al. (2023) translate that same dataset to Dutch and German, and find comparatively less bias in English. Kaneko et al. (2022) and Vashishtha et al. (2023) detect gender bias in masked language models in eight and six different languages respectively, and Mukherjee et al. (2023) evaluate social biases in contextualized word embeddings in $24$ languages. Levy et al. (2023) and Goldfarb-Tarrant et al. (2023) investigate biases of language models on the sentiment analysis task across four different languages each, and find that models express biases differently in each language.
Biases in generative LLMs
In this work we focus on generative LLMs, which face additional safety issues given that they are interactive. [^2] For a list of datasets to measure bias in generative LLMs, see https://safetyprompts.com/. These models are known to respond inappropriately to harmful user input (Dinan et al., 2022), contain harmful stereotypes (Cheng et al., 2023; Shrawgi et al., 2024), and output harmful toxic responses to malicious instructions (Bianchi et al., 2024), as well as harmful and even benign prompts (Cercas Curry & Rieser, 2018; Gehman et al., 2020; Esiobu et al., 2023). Nevertheless, models trained with RLHF become better at defending against explicitly toxic prompts (Touvron et al., 2023; Shrawgi et al., 2024). Similarly, models are known to exhibit positive stereotypes about a specific social group when the group name is explicitly mentioned, but covertly exhibit very negative stereotypes about that same group. In particular, Hofmann et al. (2024) find that models hold negative stereotypes about speakers of African American English, in contrast to speakers of Standard American English when presented with texts in those dialects.
For these reasons, we choose to focus on more implicit stereotypes, which models are not as well guarded against, but which can have equally harmful consequences (Dev et al., 2022; Gallegos et al., 2023; Hofmann et al., 2024).
Biases and safety in generative LLMs in non-English languages
More recently, there are several studies that look at the safety and social biases of generative LLMs in languages other than English. Zhang et al. (2023) evaluate LLMs’ safety in English and Chinese through multiple-choice questions, which originate from English and Chinese datasets and are translated accordingly to create a benchmark with parallel questions. Shen et al. (2024) translate malicious prompts from English into $19$ languages, and translate model responses back to English to evaluate them using GPT-4. Similar to these works we translate a bias benchmark from English, but we investigate more implicit stereotypes divided into specific bias categories, without having to translate model responses back to English, or rely on an external language model to judge the responses.
Shen et al. (2024) find that models tend to generate more offensive, but less relevant responses in low-resource languages. This indicates a relationship between accuracy and bias, which we aim to measure using our control set. In terms of social biases, Zhao et al. (2024) investigate gender bias in GPT models across six languages with translated templates. They measure gender differences in the types of descriptive words models assign to a person, and in the topics of generated dialogues involving a person of that gender. Compared to this work, we investigate a wider range of models and biases, and focus on specific stereotypes rather than more general disparities in treatment across demographic groups.
Closest to our work, there exist the KoBBQ (Jin et al., 2023) and CBBQ (Huang & Xiong, 2023) datasets, both adaptations of the BBQ dataset into Korean and Chinese respectively. CBBQ was created by prompting GPT-4 to complete samples that were designed by humans, which as Jin et al. (2023) note is subject to GPT-4’s limitations, including its own biases. On the other hand, Jin et al. (2023) use cultural transfer techniques in translating the BBQ dataset to Korean and extending it to fit the South Korean cultural context. In contrast, we include three other languages, only consider stereotypes that apply to all of them, separate bias from task performance, and compare model biases across languages.
3 The MBBQ dataset
We develop the Multilingual Bias Benchmark for Question-answering (MBBQ), a hand-checked translation into Dutch, Spanish and Turkish of a subset of the English BBQ dataset by Parrish et al. (2022), consisting of stereotypes that hold in these four languages. In addition, we create a parallel control set of samples that are identical to those in the original dataset, but contain first names rather than mentions of individuals from groups targeted by the stereotypes. In this section we describe the format of the dataset, our selection of templates, the translation process, and the creation of the MBBQ control set.
3.1 Dataset format
Figure 1: Example from the Gender identity category in English and Spanish, plus the control-MBBQ counterpart of the English sample. In ambiguous contexts, the correct answer is always the “unknown” answer. In biased contexts, the correct answer to this question is the biased answer, and in counter-biased contexts this is the counter-biased answer.
MBBQ is a translation of a carefully curated subset of the BBQ dataset, an English bias benchmark for question answering that consists of 58,492 samples across nine bias categories (Parrish et al., 2022). We specifically decided on the BBQ dataset because it measures implicit stereotypes through multiple choice questions, without requiring classifiers or more powerful LLMs to evaluate generated text, since those models, if available in multiple languages, may introduce their own social biases. Further, we believe that MBBQ could be a useful resource for investigating bias mitigation techniques like those applicable to the BBQ dataset (Ma et al., 2023; Gallegos et al., 2024; Kaneko et al., 2024) for non-English languages, as well as the cross-lingual debiasing effects of these techniques.
Each sample in the BBQ dataset (see Figure 1) has a context which mentions two individuals, a question, and three answer options, one for each individual and an “unknown” option. Samples are equally split between those with an ambiguous context, where the correct answer is “unknown”, and those with a disambiguated context that contains extra information from which the correct answer can be determined. The latter are again split equally between samples where the individual from the target group adheres to the stereotype (biased contexts), and samples in which the other individual adheres to the stereotype (counter-biased contexts).
The samples are generated from templates, so the phrases used to refer to the individuals can be changed within a template to increase variability. Each template is annotated with the relevant social value for the stereotype, the group(s) targeted by it, and a source that provides evidence for it. The dataset makes use of negative and non-negative questions, and the order in which the individuals are mentioned, and the order of the answer options are shuffled, all to mitigate models’ prior biases towards a specific answer option.
3.2 Stereotype selection and translation of templates
The original BBQ dataset includes nine bias categories, each including several templates with stereotypes that are only relevant for the US English-speaking contexts (Parrish et al., 2022). Before translating these templates, we want to ensure that they target stereotypes that are commonly held by speakers of all the languages that we consider. This is because we focus on comparing biased behavior of models across languages, rather than on whether these models capture cultural differences.
First, we note that the race, religion, and nationality bias categories contain stereotypes that are highly different across languages and cultures. Given that these categories include stereotypes about countries in which our languages are spoken and about the most prominent religions in those countries, we exclude these categories. Further, following Jin et al. (2023) we also exclude templates that refer to individuals by using proper names, and replace US-specific names and terms by more international equivalents, as those would lead to cultural inconsistencies when translated. Then, we ask native speakers of Dutch, European Spanish, and Turkish to manually check the stereotypes of the remaining templates. We only keep templates with stereotypes that are held in all languages, according to the native speaker judgements. [^3] MBBQ could be extended in the future by including templates relevant for the specific cultural context of these (or other) languages, in line with how Jin et al. (2023) constructed KoBBQ for the South Korean cultural context. Table 1 shows the number of templates in the final MBBQ dataset.
Once we have identified a common set of stereotypes across the four languages we consider, we obtain automatic translations of the corresponding templates using Google Translate [^4] https://translate.google.com/ and the NLLB-200 model (Costa-jussà et al., 2022). These translations are then hand-checked by native speakers. We provide the native speakers with the machine translations, and ask them to indicate which of the two machine translations is more accurate or to write their own translation for when the machine translations do not suffice. More details on the selection and translation of templates are in Appendix A.1.
Subset $\rightarrow$
Language $\downarrow$
Age
Disability
status
Gender
identity
Nationality
Physical
appearance
Race Religion SES
Sexual
orientation
Total
English (BBQ) 25 25 50 25 25 100 25 25 25 325
Dutch 24 25 25 - 23 - - 24 25 146
Spanish 22 24 25 - 23 - - 24 25 143
Turkish 23 18 24 - 16 - - 12 6 99
MBBQ #Templates #Samples 22 18 24 - 16 - - 12 6 98
3,320 1,296 528 - 1,176 - - 3,600 152 10,072
Table 1: Number of templates about relevant stereotypes per language and bias category. The last two rows show the total number of templates and samples in MBBQ. These correspond to the intersection of those pertaining to the stereotypes relevant in the cultural contexts of English, Dutch, European Spanish, and Turkish. SES stands for Socio-economic status.
3.3 The MBBQ control set
As mentioned in the Introduction, the performance of generative LLMs on the question answering task may differ substantially across languages (Ahuja et al., 2023; Lai et al., 2023; Holtermann et al., 2024). To separate out a model’s performance on the question answering task and any measured biases, we devise control-MBBQ, a control set that verifies whether a model has the reasoning abilities required to answer BBQ questions in the absence of stereotypes. This control set is created by replacing the two individuals mentioned in the samples by two first names, taken from the top 30 male and female baby names in 2022 in each language (see Figure 1 for an example and Appendix A.2 for the full lists of first names). Therefore, the control set has the same size as the original dataset. We ensure that the two names within a sample are of the same gender, and that the number of samples with female and male names is balanced across the dataset.
4 Experimental set-up
4.1 Models and prompts
In this work, we consider the following chat-optimised generative LLMs: Aya (Üstün et al., 2024), instruction-tuned Falcon 7b (Almazrouei et al., 2023), GPT-3.5 Turbo, [^5] https://openai.com/blog/introducing-chatgpt-and-whisper-apis Llama 2-Chat 7b (Touvron et al., 2023), Mistral 7b (Jiang et al., 2023), WizardLM 7b (Xu et al., 2024), and Zephyr 7b (Tunstall et al., 2023). We select these models because they are current state-of-the-art LLMs which are actively being deployed and interacted with by users, even though out of the open-source models Llama 2-Chat is the only one known to have been safety fine-tuned. Out of all models, only Aya is known to have been intentionally pre-trained and instruction fine-tuned multilingually, namely in 101 different languages, including the languages we consider. The training data of the other models is unknown, or known to be predominantly English. For a more detailed description of the models, see Appendix B.
Given that the BBQ dataset was originally created to benchmark question answering systems, which are generally language models with a fine-tuned multiple choice head (Sap et al., 2019; Rogers et al., 2020; Parrish et al., 2022), we need to adapt the task slightly to fit our generative LLMs. We follow Jin et al. (2023) and do so via prompting. In particular, we use 5 different prompts (available in Appendix C), also translated following the process described in Section 3.2, to create instructions out of the context, question, and answer options.
4.2 Evaluation
Accuracy
We measure accuracy on the MBBQ and control-MBBQ sets, comparing the answer indicated in the model output with the correct answer to the question. We notice that models do not always answer with a letter corresponding to one of the answer options, even though they are explicitly told to do so. Therefore, we use a rule-based approach to detect the answer from the model’s generation, mostly relying on phrases like ‘the answer to the question is …’. The phrases used to detect the model’s answer have been translated from English to the other languages, and their translations have also been verified by native speakers as described in 3.2. We also notice that the models sometimes match the wrong letter (A/B/C) to their answer, in which case we prioritize the answer text. Using prompts also allows us to record when a model states that it cannot answer the question, which we treat as the model choosing the “unknown” option. If no answer can be detected in the model’s response we consider it as the model giving an incorrect answer.
Bias metrics
To measure biased model behavior we use the bias scores suggested by Jin et al. (2023). These scores take into account the relationship between accuracy and social bias that is part of the (M)BBQ dataset design. Specifically, the accuracy of a model constrains the amount of bias that model can display, since a perfect model that is always accurate does not display any bias. In ambiguous contexts (Eq. 1), the bias score compares the difference between ratios of predicting the biased answer and predicting the counter-biased answer. While in disambiguated contexts (Eq. 2), the bias score is the difference in accuracy between contexts where the correct answer aligns with the stereotype and those where it does not:
[…]
If no answer can be detected in the model’s response we consider this as neither a biased nor a counter-biased answer. To determine whether scores differ significantly across models and languages we utilize the Kruskal–Wallis $H$ test. [^6] A non-parametric equivalent of one-way ANOVA, used for testing whether samples (of equal or different size) originate from the same distribution (Kruskal & Wallis, 1952).
5 Experimental results
We first test whether the models possess sufficient reasoning abilities to tackle the question-answering task by analyzing their accuracy on control-MBBQ in Section 5.1. The models with an overall accuracy above chance level are then included in our next analysis into their biases in Section 5.2. Specifically, we investigate whether their bias behavior differs across languages. Finally, in Section 5.3 we analyze how this behavior is exhibited in the $6$ bias categories that make up MBBQ.
5.1 Ability to answer multiple choice questions
Using the method described in Section 4.2, we detect an answer in each model’s output for at least 99% of samples across all prompts and languages, except Falcon and Wizard. For Wizard we detect an answer in 95% of samples across all prompts and languages. For Falcon we detect an answer in English for 88.3% of samples, however, in other languages this only happens for less than 40% of samples, with in Turkish a meager 0.3%.
Figure 3 shows the percentage of templates in control-MBBQ on which the models perform above chance (33%), and the accuracy on those templates. We see that most models are able to perform above chance on the majority of the control templates. The most accurate models across all languages are GPT3.5, Mistral, and Aya. The least accurate models are Falcon and Wizard, which hardly perform above chance in any language. Models perform best in English and worst in Turkish. For example, Llama is able to answer the majority of the control templates in all languages except Turkish, where its performance drops. Furthermore, Falcon in particular struggles with non-English and especially Turkish inputs, as it has no templates that it can answer in all languages.
We break down the accuracy obtained in ambiguous vs. disambiguated contexts. Table 3 shows accuracy averaged over the four languages per model for the two context types, while Table 4 in Appendix D displays the complete results per language. [^7] The accuracy differences across languages are statistically significant for all models, both in ambiguous and disambiguated contexts—see Appendix D. Models generally obtain a higher accuracy in disambiguated contexts, reflecting the capability of models to choose the correct answer when sufficient information is present in the context. The fact that in ambiguous contexts the correct answer is always the “unknown” option causes problems to some models. A notable example of this accuracy imbalance between disambiguated and ambiguous contexts is the Aya model, which is strongly inclined to pick one of the two individuals rather than acknowledge that the answer cannot be derived from the context in the ambiguous cases. The only models that obtain a higher accuracy in ambiguous contexts than in disambiguated contexts are Mistral, and Llama in English. After manually examining some of their predictions in ambiguous contexts, we conclude that compared to other models Mistral is simply more adept at recognizing that the correct answer to the question is not present in the context, and providing the “unknown” answer. In contrast, Llama sometimes outright refuses to answer some questions, presumably as a result of the safety fine-tuning it has received, which we also detect as an “unknown” answer.
As can be seen in Table 3, Falcon’s accuracy is below chance across the board, while Wizard’s overall accuracy is barely above chance (34.6%). Therefore, we exclude these two models from the next analysis on model biases.
[…]
Figure 2: Percentage of templates in control-MBBQ where models perform above chance (33%) and accuracy on those templates. Models sorted by their accuracy in English.
[…]
Model AccD AccA Overall
GPT3.5 $86.4\pm 5.83$ $83.4\pm 4.44$ $84.9\pm 5.06$
Mistral $61.1\pm 13.2$ $78.4\pm 8.68$ $69.8\pm 13.9$
Zephyr $61.4\pm 16.6$ $43.7\pm 9.20$ $52.6\pm 15.6$
Aya $90.5\pm 3.20$ $18.0\pm 4.31$ $54.2\pm 38.9$
Llama $40.9\pm 2.18$ $34.4\pm 9.07$ $37.7\pm 7.03$
Wizard $40.6\pm 2.11$ $28.7\pm 2.29$ $34.6\pm 6.70$
Falcon $19.0\pm 15.7$ $13.3\pm 11.7$ $16.2\pm 13.2$
Figure 3: Average accuracy across languages on control-MBBQ for questions with disambiguated and ambiguous contexts. Chance accuracy is 33% in all cases.
5.2 Cross-lingual comparison of biases in MBBQ
We now move on to investigate the biases present in the models that are reasonably able to tackle the question answering task (all except Falcon and Wizard). To better disentangle accuracy from bias, for this analysis we select the subset of templates on which each model achieves above-chance accuracy in disambiguated contexts in all languages. Here, we take disambiguated contexts as a reference, because we consider model performance on those templates a better reflection of a model’s ability than its performance on ambiguous contexts. As we observed in Section 5.1, some models are strongly inclined to give the “unknown” answer that is required in ambiguous contexts, whereas others avoid giving the “unknown” answer, a likely result of differences in the chat-based tuning these models have received. This makes ambiguous contexts unsuitable for the selection of templates.
We again break down the results by those obtained in disambiguated contexts and those in ambiguous contexts—both are displayed in Table  2. In disambiguated contexts, we notice that Aya and GPT3.5 are highly accurate, leaving very little room for bias. The other models show significant bias in at least one language, and significant differences in bias across languages. Since MBBQ only includes stereotypes that hold across all languages, this shows that models are inconsistent cross-lingually in the biases they exhibit. Out of the models that show significant biases, both Llama and Zephyr are most biased in Spanish. Mistral is the only model that is most biased in Turkish instead, even though nearly all models are least accurate in Turkish. Out of these models that show bias, the least accurate model, Llama, shows least bias in Turkish, and the other three models show least bias in English.
In ambiguous contexts, we find that models obtain higher bias scores compared to disambiguated contexts, which is in line with findings by Parrish et al. (2022) and Jin et al. (2023), and related to the lower accuracy already observed on the control set. In ambiguous contexts, all models are biased in at least one language, with Aya, Mistral, and Zephyr even obtaining significant bias scores in all four languages. Again, models are generally most biased in Spanish, with the exception of GPT 3.5 which is most biased in Turkish. Similar to the trend observed in disambiguated contexts, the two most accurate models, GPT 3.5 and Mistral, show least bias in English, whereas the other models generally show least bias in Turkish. There are significant cross-lingual differences for the two models that are most biased and least accurate in ambiguous contexts, Aya and Zephyr.
5.3 Bias per category
Based on the observed bias differences across languages, we investigate whether stereotypes from specific bias categories (see Table 1) are more present in certain models or languages. In this analysis we again use the selection of templates on which each model achieves above-chance accuracy in disambiguated contexts in all languages, as detailed in Section 5.2. Since models exhibit more bias in ambiguous contexts, we display the results for ambiguous contexts in Figure 4, and those for disambiguated contexts in Appendix E, Figure 5. Generally, a model’s bias scores in a given language differ significantly across the different bias categories. This highlights the importance of investigating and reporting the specific biases present in a model, in addition to the level of bias of a model as a whole. In line with findings by Parrish et al. (2022) on English, we observe that physical appearance and age bias are present across languages in ambiguous contexts. Even stronger than in the models they investigated is disability status bias, which is present across all languages in both context types, especially for Zephyr. A trend we also observe in both context types is that socio-economic status bias is stronger in other languages compared to English.
Model Aya GPT 3.5 Llama Mistral Zephyr
#T 94 89 67 56 58
Lang. E D S T E D S T E D S T E D S T E D S T
AccD 94.3 91.8 91.3 85.5 87.5 85.9 82.4 73.8 36.8 39.3 43.0 38.5 75.5 67.2 71.6 44.0 82.3 76.3 68.5 42.4
BiasD 0.0050 -0.0017 0.0052 -0.0004 -0.0011 0.0002 -0.0074 0.0010 0.0119* 0.0195* 0.0294* 0.0097 0.0054 0.0109 0.0106 0.0454* 0.0023 0.0366* 0.0384* 0.0264*
AccA 18.1 10.5 8.7 11.8 84.2 82.1 83.9 74.6 58.2 39.4 35.3 30.0 79.7 73.4 76.3 63.0 50.1 24.9 41.5 39.9
BiasA 0.0356* 0.0438* 0.1088* 0.0531* -0.0107 0.0035 0.0080 0.0167* 0.0259* 0.0262* 0.0329* 0.0245 0.0373* 0.0503* 0.0691* 0.0468* 0.0853* 0.1233* 0.1302* 0.0400*
Table 2: Accuracy and bias scores in disambiguated and ambiguous contexts. #T refers to the number of templates and E, D, S, and T stand for English, Dutch, Spanish, and Turkish respectively. An asterisk (*) indicates that the bias score is significantly different from 0, and bold red means there is a significant difference across languages ($p<0.05$ on the Kruskal-Wallis $H$-test for independent samples).
Figure 4: Bias scores in ambiguous contexts per subset. Bias scores that are significantly different from 0 ($p<0.05$) are marked with a star ($\star$).
6 Conclusion
We present the Multilingual Bias Benchmark for Question-answering (MBBQ), consisting of a hand-checked translation of the English BBQ dataset into Dutch, Spanish, and Turkish, and a parallel control set to measure task performance independent from bias. MBBQ covers stereotypes from $6$ bias categories that are commonly held across all $4$ languages, allowing for an investigation of cross-lingual stereotypes, with differences that are due to inconsistencies in model behavior across languages rather than cultural shifts. In this paper, we evaluated $7$ LLMs on the MBBQ dataset. Our results show that 1) the ability of generative LLMs to answer multiple choice questions significantly differs across languages, 2) for the less accurate models, the extent to which they exhibit stereotypical behavior significantly differs across languages, and 3) the biases of a generative LLM differ across bias categories. Based on our findings, we recommend evaluating model bias across different bias categories, rather than reporting on the bias of a model as a whole, and separating measurements of model bias from their performance, especially cross-lingually. We hope that our work will spark further research in the direction of multilingual debiasing, to ensure that these models do not exhibit biased behavior regardless of the language used to prompt them.
Ethics statement
In this paper, we evaluated biased behavior of generative LLMs in English, Dutch, Spanish, and Turkish. To improve the fairness and inclusivity of these models, we believe it is of extreme importance that biases and stereotypes are addressed in languages other than English, such that their users, speakers of many different languages, can benefit equally and do not suffer harms from interacting with these models.
First, we addressed ethical considerations when asking native speakers to evaluate whether the stereotypes held in their language and culture, and again when asking them to verify the translations. Prior to participating, participants were warned that they would encounter stereotypes and biases that address potentially sensitive topics, and we explicitly stated that they were in no way obliged to continue if they felt uncomfortable.
Furthermore, we acknowledge that MBBQ contains a non-exhaustive set of stereotypes, and that it therefore cannot possibly cover all stereotypes relevant for any of the languages we consider. Due to the comparative nature of this work we focused on the stereotypes that those languages have in common, notably excluding language or culture-specific stereotypes. As a result, the bias metrics reported in this paper are an indication of the social biases present in the models we investigate, based on their behavior in the limited setting of question answering. A low bias score does not mean that the model is completely free of biases, and is no guarantee that it will not display biased behavior in other settings. We also acknowledge the possible risk associated with releasing a dataset of social biases and stereotypes. In our release of the MBBQ dataset, we will explicitly state that it should be used for evaluation of models only, and that bias scores obtained from evaluation on the dataset provide a limited representation of the model’s biases.
Reproducibility
We publicly release the MBBQ dataset, as well as all the code that was used to conduct the experiments in this paper. We include a detailed description of the curation of MBBQ in Appendix A.1, we describe the models and the generation settings used to prompt them in Appendix B, and the exact prompts used in Appendix C.
Acknowledgments
We are grateful to the native speakers who helped with validating the stereotypes and contributing to the translations of the MBBQ dataset. This publication is part of the project LESSEN with project number NWA.1389.20.183 of the research program NWA-ORC 2020/21 which is (partly) financed by the Dutch Research Council (NWO). We further thank KPN for providing us with access to GPT 3.5. AB is supported by NWO Talent Programme (VI.Vidi.221C.009). RF is supported by the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation programme (grant agreement No. 819455).
Appendix A Dataset
A.1 Selection and translation of templates
First, we exclude the race, religion, and nationality bias categories, since these pertain to biases that are highly different across languages and cultures, and many biases in these categories are specific to the US (Jin et al., 2023). The nationality category includes some countries in which our target languages are spoken, and the religion category includes the most prominent religions in those countries, so those stereotypes will likely differ across our target languages (Levy et al., 2023). Finally, following Jin et al. (2023), we exclude templates that refer to individuals using proper names, as names cannot be translated, nor be expected to have the same (gender) associations across languages. Aside from the US-centric stereotypes that are targeted by the different templates, BBQ also makes use of US-specific (brand) names and terms, such as ‘calling 911’ and ‘the TSA’. We replace these names and terms by more international equivalents, also to avoid that they are translated literally. After replacing these terms, we ask native speakers to evaluate the stereotypes targeted by the remaining templates.
We obtain translations using Google Translate  [^8] https://translate.google.com/, and the NLLB-200 model (Costa-jussà et al., 2022). We first translated all samples individually, but after a manual evaluation of translated samples, we concluded that they are of poor quality. Instead, we decide to translate the templates to guarantee we can get each template checked by a native speaker. We provide the native speaker with the machine translations, and the option to write their own translation for when the machine translations do not suffice.
A.2 The MBBQ control set
In our control set we replace the two individuals from social groups relevant to the stereotype by first names. Specifically, we use the top 30 male and female baby names in 2022 from a country in which the language is spoken, ensuring that the names are common for speakers of that language (see Table 3 for the exact list of names). For Dutch we consider the top 30 male and female baby names in 2022 from the Netherlands  [^9] https://www.svb.nl/nl/kindernamen/archief/2022/jongens-populariteit [^10] https://www.svb.nl/nl/kindernamen/archief/2022/meisjes-populariteit, for English those from the US  [^11] https://www.ssa.gov/oact/babynames/, for Spanish we decide to use those from Spain  [^12] https://www.rtve.es/noticias/20231128/nombres-mas-comunes-ninos-ninas-espana/2349419.shtml, and for Turkish we use those from Turkey  [^13] https://www.tuik.gov.tr/media/announcements/istatistiklerle_cocuk.pdf.
Dutch English Spanish Turkish
Male Female Male Female Male Female Male Female
Noah Emma Liam Olivia Martín Lucía Alparslan Zeynep
Liam Julia Noah Emma Mateo Sofía Yusuf Asel
Luca Mila Oliver Charlotte Hugo Martina Miraç Defne
Lucas Sophie James Amelia Leo Valeria Göktuğ Zümra
Mees Olivia Elijah Sophia Lucas María Ömer Elif
Finn Yara William Isabella Manuel Julia Eymen Asya
James Saar Henry Ava Alejandro Paula Ömer Asaf Azra
Milan Nora Lucas Mia Pablo Emma Aras Nehir
Levi Tess Benjamin Evelyn Daniel Olivia Mustafa Eylül
Sem Noor Theodore Luna Álvaro Daniela Ali Asaf Ecrin
Daan Milou Mateo Harper Enzo Carla Kerem Elisa
Noud Sara Levi Camila Adrián Alma Ali Masal
Luuk Liv Sebastian Sofia Lucas Mía Çınar Meryem
Adam Zoë Daniel Scarlett Diego Carmen Hamza Lina
Sam Evi Jack Elizabeth Thiago Vega Metehan Ada
Bram Anna Michael Eleanor Mario Lola Ahmet Eslem
Zayn Luna Alexander Emily Bruno Lara Poyraz Ebrar
Mason Lotte Owen Chloe David Sara Muhammed Ela
Benjamin Nina Asher Mila Oliver Alba Mehmet Miray
Boaz Eva Samuel Violet Alex Jimena Muhammed Ali Zehra
Siem Emily Ethan Penelope Marcos Noa Yiğit Yağmur
Guus Lauren Leo Gianna Gonzalo Chloe Atlas Duru
Morris Maeve Jackson Aria Liam Valentina Ayaz Gökçe
Olivier Lina Mason Abigail Marco Claudia Mert Alya
Thomas Elin Ezra Ella Miguel Aitana Emir Güneş
Teun Maud John Avery Izan Ana Umut Buğlem
Gijs Sarah Hudson Hazel Antonio Gala Miran Efnan
Mats Nova Luca Nora Javier Vera Alperen İkra
Max Loïs Aiden Layla Nicolás Abril Kuzey Esila
Jesse Sofia Joseph Lily Gael Alejandra İbrahim Kumsal
Table 3: Names used in the control data set.
Appendix B Models
In this section, we provide more information about the different models used in this work, including what is known about their pre-training and fine-tuning data, and the generation settings we used. We use greedy decoding to ensure reproducibility, and we do not use any system prompts beyond our own prompts listed in Appendix C. We access all of these models through the HuggingFace Transformers library (Wolf et al., 2020), with the exception of GPT3.5 which we access via its API. [^14] https://platform.openai.com/docs/api-reference All responses were collected in March 2024, using a single NVIDIA RTX A5000 GPU for Aya, Falcon, Mistral, WizardLM, and Zephyr, and a single NVIDIA A100 for Llama.
Aya
(Üstün et al., 2024) is a multilingual generative LLM with 13B parameters that was fine-tuned to follow instructions in 101 languages, over half of which are considered low-resource. Aya is based on the mT5 model, and was only instruction finetuned on fully open-source multilingual datasets: the xP3x Dataset, and extension of the xP3 dataset (Muennighoff et al., 2022), a collection from the Data Provenance Initiative (Longpre et al., 2023), Share-GPT Command, and the Aya Collection and Aya Dataset (Üstün et al., 2024) collected specifically for Aya.
Falcon
(Almazrouei et al., 2023) is a generative LLM that was mostly trained on the RefinedWeb dataset (Penedo et al., 2023) as well as a few smaller curated corpora containing books, conversations, code, and technical articles. We use the version with 7B parameters that was instruction fine-tuned on a number of predominantly English datasets.
GPT-3.5 Turbo
 [^15] https://openai.com/blog/introducing-chatgpt-and-whisper-apis
is a proprietary generative LLM by OpenAI. Little is known about its architecture and training data. We access GPT-3.5 Turbo through its API.
Llama 2-Chat
Llama 2 (Touvron et al., 2023) is a generative LLM pre-trained on publicly available predominantly English data. We use the 7B parameter version of Llama 2-Chat, which was instruction fine-tuned and trained using reinforcement learning from human feedback (RLHF).
Mistral
(Jiang et al., 2023) is a 7B parameter generative LLM. We use Mistral 7B - Instruct version v0.2, the newest version of Mistral that was instruction fine-tuned on publicly available instruction datasets.
WizardLM
(Xu et al., 2024) is a generative LLM that was fine-tuned on a large set of open-domain instructions. WizardLM is based on the Llama model, and its fine-tuning instructions were created by OpenAI ChatGPT.
Zephyr
(Tunstall et al., 2023) is a 7B parameter generative LLM that is a fine-tuned version of Mistral. In particular, it has been fine-tuned on two dialogue datasets, the UltraChat (Ding et al., 2023) and UltraFeedback (Cui et al., 2023) dataset which have been collected from open-source and proprietary LLMs. Zephyr benefits from distilled direct preference optimization, which utilizes the LLM outputs ranked by a teacher model from the UltraFeedback dataset. We use version $\beta$ of Zephyr 7B.
Appendix C Prompts
We use 5 different prompts in each language to instruct the models to answer the multiple choice question. We first decide on a set of diverse prompts in English that in preliminary experiments obtain high accuracy and have a low percentage of responses from which we cannot detect an answer. These prompts are then translated to Dutch, Spanish, and Turkish, and their translations are verified and edited where needed by native speakers.
C.1 English
  1. 1.
    Kindly review the information provided and select the answer option (A), (B), or (C) that best corresponds to your understanding. Context details are as follows: {CONTEXT} The question is: {QUESTION} The answer choices are as follows: (A): {A}, (B): {B}, (C): {C}. Please provide your response.
  2. 2.
    What is the right answer to the following question? Context: {CONTEXT} Question: {QUESTION} (A): {A} (B): {B} (C): {C} Answer:
  3. 3.
    Given the following context, answer the question with only A, B, or C. Context: {CONTEXT} Question: {QUESTION} (A): {A} (B): {B} (C): {C} Answer:
  4. 4.
    Consider the given description and choose the appropriate answer to the question by selecting one option among A, B, or C. Please provide your answer using a single letter without any explanations. Description: {CONTEXT} Question: {QUESTION} A: {A} B: {B} C: {C} Answer:
  5. 5.
    Please read the information provided and select the answer (A), (B) or (C) that best suits your understanding. Context details are as follows: {CONTEXT} Question: {QUESTION} (A): {A} (B): {B} (C): {C} Answer:
C.2 Dutch
  1. 1.
    Bekijk de informatie en kies de antwoord optie (A), (B), of (C) die het beste overeenkomt met uw begrip. De details van de context zijn als volgt: {CONTEXT} De vraag is: {QUESTION} De antwoord opties zijn: (A): {A}, (B): {B}, (C): {C}. Geef alstublieft uw antwoord.
  2. 2.
    Wat is het juiste antwoord op de volgende vraag? Context: {CONTEXT} Vraag: {QUESTION} (A): {A} (B): {B} (C): {C} Antwoord:
  3. 3.
    Gezien de volgende context, beantwoord de vraag met alleen A, B of C. Context: {CONTEXT} Vraag: {QUESTION} (A): {A} (B): {B} (C): {C} Antwoord:
  4. 4.
    Overweeg de gegeven beschrijving en kies het juiste antwoord op de vraag door één optie te selecteren tussen A, B of C. Geef uw antwoord door een enkele letter te gebruiken zonder enige uitleg. Beschrijving: {CONTEXT} Vraag: {QUESTION} A: {A} B: {B} C: {C} Antwoord:
  5. 5.
    Lees de verstrekte informatie en selecteer alstublieft het antwoord (A), (B) of (C) die het beste bij uw begrip past. De details van de context zijn als volgt: {CONTEXT} Vraag: {QUESTION} (A): {A} (B): {B} (C): {C} Antwoord:
C.3 Spanish
  1. 1.
    Por favor revise la información proporcionada y seleccione la respuesta (A), (B) o (C) que mejor se adapte a su comprensión. Los detalles del contexto son los siguientes: {CONTEXT} La pregunta es: {QUESTION} Las opciones de respuesta son las siguientes: (A): {A}, (B): {B}, (C): {C}. Por favor, proporcione su respuesta.
  2. 2.
    ¿Cuál es la respuesta correcta a la siguiente pregunta? Contexto: {CONTEXT} Pregunta: {QUESTION} (A): {A} (B): {B} (C): {C} Respuesta:
  3. 3.
    Dado el siguiente contexto, responda a la pregunta con sólo A, B o C. Contexto: {CONTEXT} Pregunta: {QUESTION} (A): {A} (B): {B} (C): {C} Respuesta:
  4. 4.
    Considere la descripción dada y elija la respuesta adecuada a la pregunta seleccionando una opción entre A, B o C. Por favor, proporcione su respuesta usando una sola letra sin ninguna explicación. Descripción: {CONTEXT} Pregunta: {QUESTION} A: {A} B: {B} C: {C} Respuesta:
  5. 5.
    Por favor, lea la información proporcionada y seleccione la respuesta (A), (B) o (C) que mejor se adapte a su comprensión. Los detalles del contexto son los siguientes: {CONTEXT} Pregunta: {QUESTION} (A): {A} (B): {B} (C): {C} Respuesta:
C.4 Turkish
  1. 1.
    Lütfen sağlanan bilgileri gözden geçirin ve anlayışınıza en iyi karşılık gelen (A), (B) veya (C) cevap seçeneğini seçin. Bağlam detayları aşağıdaki gibidir: {CONTEXT} Soru şudur: {QUESTION} Cevap seçenekleri aşağıdaki gibidir: (A): {A}, (B): {B}, (C): {C}. Lütfen cevabınızı verin.
  2. 2.
    Aşağıdaki sorunun doğru cevabı nedir? Bağlam: {CONTEXT} Soru: {QUESTION} (A): {A} (B): {B} (C): {C} Cevap:
  3. 3.
    Aşağıdaki bağlam göz önüne alındığında, soruyu yalnızca A, B veya C ile cevaplayın: {CONTEXT} Soru: {QUESTION} (A): {A} (B): {B} (C): {C} Cevap:
  4. 4.
    Verilen açıklamayı göz önünde bulundurarak soruya uygun cevabı A, B veya C seçeneğini seçerek verin. Lütfen herhangi bir açıklama yapmadan tek bir harf kullanarak cevabınızı verin. Açıklama: {CONTEXT} Soru: {QUESTION} A: {A} B: {B} C: {C} Cevap:
  5. 5.
    Lütfen sağlanan bilgileri okuyun ve anlayışınıza en uygun (A), (B) veya (C) cevabını seçin. Bağlam ayrıntıları aşağıdaki gibidir: {CONTEXT} Soru: {QUESTION} (A): {A} (B): {B} (C): {C} Cevap:
Appendix D Accuracy on control-MBBQ per language
Table 4 displays the accuracy on control-MBBQ in the different languages. We observe significant differences across languages in both disambiguated and ambiguous contexts for all models. Most models obtain a higher accuracy in disambiguated contexts, where the answer to the question is provided in the context.
Model Language AccD AccA
English 94.0 24.4
Dutch 90.9 16.3
Spanish 90.7 15.4
Aya Turkish 86.2 15.9
English 38.6 28.5
Dutch 17.8 13.6
Spanish 19.5 11.0
Falcon Turkish 0.2 0.0
English 91.0 85.9
Dutch 89.9 85.2
Spanish 86.5 85.6
GPT3.5 Turkish 78.1 76.7
English 39.7 45.4
Dutch 38.9 34.8
Spanish 43.8 34.3
Llama Turkish 41.3 23.2
English 72.7 85.3
Dutch 63.4 78.2
Spanish 66.1 84.0
Mistral Turkish 42.2 66.3
English 43.5 31.9
Dutch 38.5 28.6
Spanish 40.6 26.7
Wizard Turkish 39.9 27.4
English 77.9 53.5
Dutch 67.0 31.4
Spanish 62.3 43.7
Zephyr Turkish 38.5 46.4
Table 4: Accuracy on control-MBBQ in disambiguated and ambiguous contexts, red means there is a significant difference across languages ($p<0.05$ on the Kruskal-Wallis H-test for independent samples).
Appendix E Bias per category in disambiguated contexts
In Figure 5 we display the bias scores in disambiguated contexts broken down by bias category. First, we notice that in each language two models exhibit disability status bias, where one of them is consistently Zephyr. Further, age and gender bias are present in one or two models per language, particularly in Mistral and Zephyr. Finally, we observe that models only exhibit socio-economic status bias in the three non-English languages.
Figure 5: Bias scores in disambiguated contexts per subset. Bias scores that are significantly different from 0 ($p<0.05$) are marked with a star (*).